### v1

# Hardware test and validation of the angular droop control: Analysis and experiments

Abstract
The angular droop control is a grid-forming control strategy that exploits the idea of power-to-angle droop to achieve exact frequency synchronization with no stringent separation between primary and secondary frequency control. In this work, we conduct hardware experiments in the Smart Energy System Control Laboratory at Karlsruhe Institute of Technology (KIT) to test and validate the angular droop control for low voltage power grids in two different test scenarios. First, we verify its grid-forming capabilities after a major event, e.g., following a blackout, demonstrated via power-to-angle droop behavior. For this, we propose two implementation schemes that rely either on direct or indirect actuation of the modulation signal and draw a comparison between them. Second, we investigate the plug-and-play capabilities, i.e., local stability and power sharing for a two-converter system and provide suitable tuning for the control gains. Our experimental findings illustrate the usefulness of hardware test and validation for DC/AC converter control, the practical challenges entailed and the proposed remedies.
I Introduction
Power grids are facing a rapid transition from fossil fuel towards an increasing share of renewable energy resources. This profound change is characterized by the integration of converter-based generation [1]. In particular, the high penetration of power electronics alters the power system dynamics, governed thus far by rotating synchronous machines underpinning the legacy grid [2]. Therefore, the control of DC/AC converters lies at the forefront of this transition to ensure the stability of power systems [3, 4, 5].
For a given control scheme, the sole deployment of numerical simulations for validation is unsatisfactory due to the discrepancy between real-world setup and simplified settings adopted in numerical case studies such as anomalous and unmodeled dynamics, erroneous model parameters and unknown disturbances affecting the grid [5]. In fact, conducting hardware experiments in controlled laboratory environments offers the possibility to integrate hardware prototypes and is thus a significant step towards field application of new control paradigms. For this, a power-hardware-in-the-loop testbed for intelligent operation and control of low-inertia power systems is proposed in [6]. Therein, the facility is specifically designed for experimentally testing and validating control schemes for low-inertia power systems. In the same vein, the Smart Energy System Control Laboratory at KIT is introduced in [7]. It consists in a fully-automated and user-oriented research infrastructure for controlling and operating smart energy systems thanks to its high level of automation and the capacity to fully function in a grid-decoupled way. This allows to test, study and evaluate control algorithms in a safe environment.
In this context, hardware-in-the-loop tests have been conducted on several controllers to study their closed-loop capabilities under realistic operating conditions. One of the most well-studied controllers in converter-based power grid is the frequency droop control initially proposed in [8]. Inspired by the dynamics governing synchronous machines and their analogy to coupled oscillator dynamics [9], frequency droop control is a grid-forming method extensively studied both in theory and practice. An experimental validation of the droop control strategy can be found in [10] for a single and two droop-controlled converters. The dynamics of synchronous machines remain a source of inspiration for a multitude of converter control strategies that emulate the behavior of synchronous machines. One particular controller that relies on exact model matching of high-order dynamics of three-phase synchronous machines with three-phase balanced and averaged DC/AC converters derived from first-order principles is the matching control introduced in [11, 12]. The particularity of this controller relies on easily measured DC-side voltage representing an indicator of power imbalance in the grid without prior assumptions on quasi-stationary steady state and the operation on phasor quantities. An experimental setup consisting of two voltage source converters for the matching control is proposed in [13] and performed on a general-purpose test bench. Oscillator-based control schemes such as the virtual oscillator control [14] rely on emulating the dynamics of weakly coupled nonlinear oscillators. Compared to frequency droop control, which is only well-defined in the vicinity of a sinusoidal steady state, virtual oscillator control enables interconnected converters to stabilize synchronous sinusoidal waveforms starting from arbitrary initial condition. The virtual oscillator control is validated experimentally in [15] within a laboratory hardware prototype to demonstrate the validity of the design approach. Even though the virtual oscillator control has provable droop properties [16], the control tuning remains a difficult task due to a lack of intuition on the physical interpretation of the gains. It was also not possible to track active and reactive power setpoints in its original formulation in [14]. These limitations have motivated a variant of virtual oscillator control suggested in [17] that allows active and reactive power to be dispatched, hence the name dispatchable Virtual Oscillator Control or d-VOC. The d-VOC consists of a combination of a synchronizing feedback term, together with a decentralized magnitude control law. It allows for global stabilization of the angles and voltage magnitudes at their desired values, corresponding to a pre-specified solution of the AC power-flow equations. The controller exhibits droop behavior around the standard operating point, which makes it backward compatible with existing power system components. Experimental validation of the dispatchable virtual oscillator on a testbed are found in [18].
The angular droop control trades linearly active power with angle deviation at steady state and thus achieves exact frequency regulation. Compared to frequency droop control, the angular droop controller exploits the idea of power-to-angle droop to achieve frequency synchronization with zero stationary frequency error and thus merging primary with secondary frequency control [19]. The angular droop control is shown to be inverse optimal stabilizing for the angle dynamics under mild assumptions [20, 21]. Thereby, the converter phase angles are stabilized to an induced steady-state, where the convergence rate is traded with the allowed control effort. The optimality of the angular droop control brings about inherent desirable gain margins analogous to linear quadratic regulators and showcases the utility of inverse optimal control theory in networked settings [22]. The angular droop control has been studied for constrained DC/AC converter angles and overall power generation [23], for discrete-time systems [24] and numerically tested on several power system benchmarks [25, 26, 27, 20]. In the present work, we study the angular droop control in a laboratory environment provided by the Smart Energy System Control Laboratory at KIT from the lenses of hardware experimentation. First, we demonstrate that a detailed high-order model including a boost converter can be accounted for even if the theoretical analysis in [20, 21] relies on a simplified DC/AC converter model description. For this, we actuate the modulation signal representing the main input to the high-dimensional DC/AC converter. Second, we test the grid-forming capabilities of the angular droop control, namely the ability to form sinusoidal voltage waveforms at a desired AC frequency and amplitude. This is exemplified through the black start capabilities following a major event, e.g., a blackout as well as the ability to withstand load disturbances, i.e., upon a sudden increase/decrease in the load power. Third, motivated by the wide use of cascaded inner loops for the control of DC/AC converters, we propose an indirect implementation scheme for the angular droop control and draw a comparison to the direct scheme acting on the modulation signal. Fourth, we validate frequency synchronization and the power sharing capabilities of the angular droop control in dependence of the control gains shown in [20]. Therein, we provide useful insights into the tuning of the gain from a practitioner point of view. Finally, we highlight the practical challenges encountered within the different experiments and provide tractable solutions to circumvent them.
The paper is organized as follows: Section II introduces the angular droop control and discusses its properties. A detailed high-order modeling of the converters and their controls are presented in Section III. Section IV describes the experimental setup. Sections V and VI summarize the results of the test scenarios and Section VII concludes the paper.
Notation: Define $\mathbf{I}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ and $\mathbf{J}=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}.$ Let $\mathbb{T}^{n}$ denote the $n-$th dimensional torus. Consider a network described by a connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\Xi)$, consisting of $|\mathcal{V}|=n$ nodes representing DC/AC converter buses and $|\mathcal{E}|=m$ edges modeling purely inductive transmission lines (i.e., with zero conductances) with susceptance $b_{kj}>0,\;(k,j)\in\mathcal{E}$ collected in the diagonal matrix $\Xi=\mathrm{diag}(b_{kj}),\;(k,j)\in\mathcal{E}$. The topology of the graph $\mathcal{G}$ is described by the incidence matrix $\mathcal{B}\in^{n\times m}$. We denote by $\mathcal{N}_{k}$ the neighbor set of converter $k$. Let $x$ denote an AC quantity in $abc-$frame and $x_{dq}:=\mathcal{P}(\theta_{dq})\,x$ denote its transformation in $dq-$frame [2] following a Park transformation $\mathcal{P}(\theta_{dq})$ with angle $\theta_{dq}(t):=\theta_{k}(t)$.
II The angular droop control
II-A Control law
Consider a network of DC/AC converters, each represented by a voltage phasor. Hereby all the phasors are modeled with a constant magnitude (e.g. one per unit), and the converter’s angle dynamics are assumed to be controllable. We assume quasi-stationary steady state operation, i.e., the frequency is around a nominal steady state $\omega^{*}$ [2]. Overall the network dynamics can be represented by,
$\dot{\theta}=\hat{u}\,(\theta)+\omega^{*}\mathds{1}_{n},\quad\theta(0)=\theta_{0},$ (1)
where $\hat{u}(\theta)=[\hat{u}_{1}(\theta),\dots,\hat{u}_{n}(\theta)]^{\top}\in\mathbb{R}^{n}$ is the main control input, $\theta=[\theta_{1},\dots,\theta_{n}]^{\top}\in\mathbb{R}^{n}$ is the vector of phase angles of the DC/AC converters and $\theta_{0}=[\theta_{0,1},\dots,\theta_{0,n}]^{\top}\in\mathbb{R}^{n}$ is the initial angle vector. By discarding the internal converter dynamics, we design a mathematical tractable approach to find an optimal controller in a concise and closed-form based on inverse optimality as shown in [20]. In Section III and IV, we demonstrate through hardware experimentation that a detailed high-order converter model can be accounted for.
Next, we denote the vector of nominal phase angles, rotating at nominal frequency $\omega^{*}>0$ by $\theta^{*}(t)=\omega^{*}\mathds{1}_{n}+\theta^{*}_{0}\in^{n}$, where $\theta^{*}_{0}=[\theta^{*}_{0,1},\dots,\theta^{*}_{0,n}]\in^{n}$ is the initial nominal angle vector. Let $P(\theta)=[P_{1}(\theta),\dots,P_{e,n}(\theta)]^{\top}\in^{n}$ be the vector of active power injected into the network and $P^{*}=[P^{*}_{1}(\theta),\dots,P^{*}_{e,n}(\theta)]^{\top}\in^{n}$ is the vector of nominal active power. The angular droop control is given by,
$\hat{u}(\theta)=-\frac{1}{2}R^{-1}\,(\Gamma\,(\theta-\theta^{*})+P(\theta)-P^{*}),$ (2)
with $R=\text{diag}(\alpha_{1},\dots,\alpha_{n})>0,\,\Gamma=\text{diag}(\gamma_{1},\dots,\gamma_{n})>0$. To design the angular droop control (2), we assume available synchrophasor measurements with respect to a global frame of reference. Our assumption is reasonable since phasor measurements are available at the Smart Energy System Control Laboratory at KIT [7]. See also Section IV. In summary, the closed-loop angle dynamics are given by,
$\dot{\theta}=-\frac{1}{2}R^{-1}\big{(}\Gamma\,(\theta-\theta^{*})+P(\theta)-P^{*})+\omega^{*}\mathds{1}_{n}.$ (3)
II-B Properties: Stability and optimality
For clarity of exposition, we introduce the error coordinates $\tilde{\theta}(t)=\theta(t)-\theta^{*}(t)$, the angle vector at induced steady state $\theta^{s}:=\lim_{t\to\infty}\theta(t)$ and consider the following optimal control problem,
$\min_{u\in^{n}}$ $\int_{0}^{\infty}\sum_{k=1}^{n}\bigg{(}\alpha_{k}u_{k}^{2}(\tilde{\theta})+\frac{1}{4\alpha_{k}}\Big{(}\gamma_{k}\tilde{\theta}_{k}+P_{k}(\tilde{\theta})-P^{*}_{k}\Big{)}^{2}\bigg{)}\,\mathrm{d}t,$
[…]
Assumption 1.
[20] The induced steady state angle vector $\tilde{\theta}^{s}=\{\tilde{\theta}^{s}_{k}\}_{k=1}^{n}$ satisfies, $\mathcal{B}^{\top}\tilde{\theta}^{s}\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right)^{m}$, where $\mathcal{B}\in^{n\times m}$ is the incidence matrix of the underlying graph $\mathcal{G}$.
Under Assumption 1, the angular droop control is the optimal stabilizing solution of (II-B) in a neighborhood of the induced steady state angle $\tilde{\theta}^{s}$ that satisfies
$\Gamma\,\tilde{\theta}^{s}=P^{*}-P^{s}(\tilde{\theta}^{s}),$ (5)
where $P^{s}(\tilde{\theta}^{s})=[P_{1}^{s}(\tilde{\theta}^{s}),\dots,P_{n}^{s}(\tilde{\theta}^{s})]^{\top}$ is vector of the induced steady state active power [20] . Note that (5) describes the steady state as a power balance between the active power and angle deviation from the nominal value. It coincides with the steady state resulting from letting the running cost in (II-B) go asymptotically to zero, i.e.,
$\lim\limits_{t\to\infty}\left(\gamma_{k}\tilde{\theta}_{k}(t)+P_{k}(\tilde{\theta}(t))-P^{*}_{k}\right)=0,\quad\forall\,k=1,\dots,n.$
The angle droop behavior (5) leads to zero stationary frequency error. In fact, by taking the time derivative of (5), we arrive at $\dot{\tilde{\theta}}^{s}=0$ and therefore $\dot{\theta}^{s}_{k}=\omega^{*}$. Thus, the steady state frequency error is zero and we conclude that:
  • The angular droop control (5) is able to guarantee both primary and secondary frequency control at once.
  • Local asymptotic stability of the induced steady state angle $\theta^{s}$ implies local frequency synchronization of all converters at the nominal value $\omega^{*}$.
II-C Control tuning
After selecting the gain matrices $R=\alpha\,\mathbf{I}_{n}$ and $\Gamma=\gamma\,\mathbf{I}_{n}$ with $\alpha,\gamma>0$, i.e., the control gains are uniform across all the converters, observe that:
  • A decrease in the gain $\alpha>0$ improves the angle transients, i.e., it results in faster convergence of the angles towards the induced steady state angle.
  • The gain $\gamma>0$ defines the power-to-angle droop behavior between the power and angle deviation at steady state.
  • Both the gains $\gamma$ and $\alpha>0$ affect the rate of change of frequency or RoCoF given by $\dot{\omega}_{k}$. This can be seen as follows. By taking the time derivative of (3), we obtain,
    $\ddot{\theta}_{k}=-\frac{1}{2\alpha}\big{(}\gamma(\dot{\theta}_{k}-\dot{\theta}^{*}_{k})+\dot{P}_{k}(\theta)\big{)},\quad k=1\dots n$
    where $\dot{P}_{k}(\theta)=\frac{\mathrm{d}P_{k}(\theta)}{\mathrm{d}t}$. By letting $\omega_{k}:=\dot{\theta}_{k}$ and $\omega^{*}_{k}:=\dot{\theta}^{*}_{k}$, we have
    $\dot{\omega}_{k}=-\frac{1}{2\alpha}\big{(}\gamma\,(\omega_{k}-\omega^{*})+\dot{P}_{k}(\theta)\big{)},\quad k=1\dots n\,.$ (6)
    Therefore, a sudden change in active power corresponds to a sudden RoCoF $\dot{\omega}_{k}$ that occurs during the transients, while the frequency error remains zero at steady state. In (6), the RoCoF depends on both gains $\alpha>0$ and $\gamma>0$.
III High-order converter modeling and control
With some abuse of notation, we omit throughout subsequent sections the subscript $k$ to denote a quantity $z_{k}$ of the $k-$th converter.
III-A Modeling
In this section, we relax our modeling assumptions from Section II towards a hardware experimentation in two directions:
  • First, we include a DC power supply and a DC/DC converter put in series behind the DC/AC converter and suggest a control approach for the boost converter relying on vector control consisting of nested voltage and current control loops.
  • Second, even though the derivation of the angular droop control ignores the internal dynamics of the converter, we present a sufficiently detailed, high-order model of the DC/AC converter. Based on it, we suggest a direct and indirect method to implement the angular droop control (2) by modulation control.
III-A1 Boost converter
Figure 1: A schematic representation of the boost converter under study
Fig. 1 depicts the boost converter in our experimental setup. Observe that the two-switch configuration is due to the available half-bridge modules. However, only the lower switch is actuated to perform the desired control action. The upper switch is actuated complementary to reduce the losses through the parallel diode. The boost converter is modeled by following ordinary differential equations,
$L_{b}\dot{I}_{b}$ $=-R_{b}\,I_{b}+V_{b}-V_{c}$ (7)
$C_{dc}\dot{V}_{dc}$ $=-G_{dc}V_{dc}+I_{dc},\quad I_{dc}=\frac{V_{b}I_{b}}{V_{dc}}\,.$
In (7), we denote by $I_{b}\in$ the current flowing out of the DC supply and by $V_{b}$ the DC supply voltage. The conductance $G_{dc}=1/R_{dc}>0$ models the parasitic losses on the DC side. The inductance is represented by $L_{b}$ and the DC capacitance is given by $C_{dc}>0$. Additionally, $V_{dc}\in$ represents the voltage across the DC capacitor and $V_{c}\in$ is the voltage controlled directly by the duty cycle $d\in[0,1]$ via the relationship
$V_{c}=(1-d)\cdot V_{dc}\in.$ (8)
Note that the duty cycle $d$ in (8) represents the main control input to the boost converter.
III-A2 DC/AC converter
Figure 2: A three-phase DC/AC converter model under study
In the remainder, we consider a three-phase averaged and balanced DC/AC converter as shown in Fig. 2 and described in $abc-$frame [2, 28],
$L\,\dot{i}$ $=-R\,i+\frac{1}{2}\overline{u}\,\,V_{dc}-v,$ (9)
$C\,\dot{v}$ $=-G\,v+i-i_{o},$
where $V_{dc}$ denotes the DC side voltage from the boost converter model (7). On the AC side, let $i\in^{3}$ denote the inductance current and $v\in^{3}$ the output voltage. The filter resistance and inductance are specified by $R>0$ and $L>0$, respectively. The capacitor $C>0$ is set in parallel with the load conductance $G>0$ to ground. The DC/AC converter is connected to the AC network, where $i_{o}\in$ is the output current flowing into the network. Note that the pulse width modulation signal $\overline{u}\in[-1,1]$ relates to the converter duty cycle $d_{AC}\in[0,1]$ via
$d_{AC}=\frac{1}{2}+\frac{\overline{u}}{2},$ (10)
where $\overline{u}$ represents the main input to the DC/AC converter.
III-B Control schemes
In the remainder, we present control schemes based on vector control for the boost converter and angular droop control for the DC/AC converter.
III-B1 Boost converter
Figure 3: Summary of the vector control of the boost converter using as main input the duty cycle $d$ in (13).
The cascaded control architecture of the boost converter exploits the differential equations (7) and is summarized in Fig. 3. In particular, an outer loop regulates the DC capacitor voltage $V_{dc}$ at a nominal value $V_{dc}^{*}>V_{b}>0$ by specifying a reference current $I^{d}_{b}\in$ given explicitly by,
$I^{d}_{b}=\!\!\!\frac{V_{dc}}{V_{b}}\cdot\bigg{(}G_{dc}V_{dc}-k_{P}(V_{dc}-V^{*}_{dc})-k_{I}\int_{0}^{t}(V_{dc}(\tau)-V^{*}_{dc})\mathrm{d}\tau\bigg{)},$ (11)
where $k_{P}>0,\,k_{I}>0$ are proportional and integral control gains. The reference current $I^{d}_{b}$ in (11) is tracked by an inner current control loop leveraging the reference voltage $V^{d}_{l}:=V_{b}-V_{c}$ as follows,
$V^{d}_{l}=R_{b}I_{b}-k_{BP}\big{(}I_{b}-I^{d}_{b}\big{)}-k_{BI}\int_{0}^{t}\big{(}I_{b}(\tau)-I^{d}_{b}\big{)}\mathrm{d}\tau,$ (12)
with $k_{BP},\,k_{BI}>0$. The choice of the gain pairs $(k_{P},k_{I})$ and $(k_{BP},k_{BI})$ ensures time-scale separation, namely that the closed-loop dynamics of the current loop is faster than that of DC voltage. Finally, the duty cycle $d$, i.e., the main input to the boost converter is deduced from the reference voltage $V^{d}_{l}$ in (8) and (12) via the relationship,
$d=1-\frac{V_{b}-V^{d}_{l}}{V_{dc}}.$ (13)
III-B2 Control of DC/AC converter
We distinguish two approaches to implement the angular droop control (2) following direct or indirect control of the modulation signal $\overline{u}$ in (9).
Direct implementation
First, we define the active power $P=v^{\top}i_{o}$, and the nominal steady state $P^{*}=v^{*\top}i_{o}^{*}$ and propose the following control scheme
$\dot{\theta}$ $=-\frac{1}{2\alpha}\left(\gamma(\theta-\theta^{*})+(P-P^{*})\right)+\omega^{*},$ (14a)
$\overline{u}$ $=A\begin{bmatrix}\sin(\theta)\\ \sin(\theta-\frac{2\pi}{3})\\ \sin(\theta+\frac{2\pi}{3})\end{bmatrix},$ (14b)
where $10<A<$ 1 is the amplitude of the modulation signal $\overline{u}$. Fig. 4 depicts a summarizing block diagram of a single DC/AC converter whose system dynamics are given by (9), set in closed-loop with the angular droop control (2). Note that in (14), the angular droop control increments the converter’s state with a virtual angle. In this scheme, the modulation signal $\bar{u}$ has the virtual angle $\theta$ given by (14).
Figure 4: Implementation of the angular droop for the DC/AC converter via direct control of the modulation signal $\overline{u}$.
Indirect Implementation
In this section, we propose an indirect implementation of the angular droop control that relies on vector control. In particular, the indirect implementation entails inner voltage and current control loop according to a cascaded architecture.
After a Park transformation $\mathcal{P}(\theta_{dq})$ with angle $\theta_{dq}(t):=\theta(t)$ and given the reference voltage $v^{d}$ in $abc$-frame [2],
$v^{d}(\theta)=V^{*}\begin{bmatrix}\sin(\theta)\\ \sin(\theta-\frac{2\pi}{3})\\ \sin(\theta+\frac{2\pi}{3})\end{bmatrix},$ (15)
with $V^{*}>0$ the reference voltage amplitude and $\theta$ the angle given by (3), the tracking of the reference voltage (15) is achieved via cascaded voltage and current loops implementing proportional-integral (PI) controllers. Thereby, the outer voltage loop generates a reference current signal
$i_{dq}^{d}$ $=Y\,v_{dq}+i_{o,dq}-k_{VP}(v_{dq}-v_{dq}^{d})$ (16)
$-k_{VI}\int_{0}^{t}\big{(}v_{dq}(\tau)-v_{dq}^{d}\big{)}\mathrm{d}\tau,$
where $k_{VP},k_{VI}~{}>0$ are the control gains and $Y=G+C\mathbf{J}\omega^{*}$. To track the reference current $i_{dq}^{d}$ in (16), we design an inner current loop based on proportional and integral control using the switching voltage $v^{d}_{m}=\frac{1}{2}\overline{u}_{dq}\,V_{dc}$ as follows
$v_{m}^{d}=Z\,i_{dq}+v_{dq}-k_{IP}(i_{dq}-i^{d}_{dq})-k_{II}\int_{0}^{t}(i_{dq}(\tau)-i_{dq}^{d})\,\mathrm{d}\tau,$
where $k_{IP},k_{II}>0$ are control gains and $Z=R+L\,\mathbf{J}\,\omega^{*}$. By applying the inverse Park transformation $\mathcal{P}^{-1}(\theta_{dq})$, we recover the modulation input $\overline{u}$ in $abc-$frame as follows,
$\overline{u}(\theta)=2\;\frac{\mathcal{P}^{-1}(\theta_{dq})(v^{d}_{m})}{V_{dc}}.$ (17)
It is noteworthy that, the pairs $(k_{VP},k_{VI})$ and $(k_{IP},k_{II})$ are again chosen to guarantee time-scale separation, where the current control loop is faster than the voltage control loop. Fig. 5 summarizes the overall scheme of the indirect implementation based on the vector control.
Figure 5: Indirect cascaded control of the angular droop after $dq-$transformation of the DC/AC converter.
IV Experimental Setup
The following hardware experiments are conducted at the Smart Energy System Control Laboratory at KIT. For a comprehensive description of the overall lab equipment, see [7].
IV-A Hardware components
The hardware testbed consists of a programmable DC/AC converter system, a resistive load and transmission line replicas.
IV-A1 Programmable DC/AC converter system
As shown in Fig. 6, a programmable DC/AC converter system consists of a DC/DC converter, a control unit, a DC/AC converter and an output filter.
  • DC/DC converter: a DC/DC stage, consisting of a boost converter, is deployed to regulate the supplied DC voltage to a desired value. In the test bench, a DC power supply provides a constant DC voltage. The half-bridge power modules are intended for rapid prototyping for power electronics [29]. The adopted control scheme is described in Sec. III-B1.
  • Control unit: the real-time control unit includes an analog front-end to connect the sensor measurements, a dual-core 1 GHz ARM processor to run user-generated code. The control unit is capable of handling fast tasks such as the gate pulse generation for power electronics and enables closed-loop control with a switching frequency up to $250$ kHz [30].
  • DC/AC converter and output filter: the two-level three-phase DC/AC converter system consists of half-bridge modules described in [29]. The output filter consists of an inductance in series with a parasitic resistance. Both are set in parallel with a capacitance [31]. The DC/AC converter system has a rated power of $P_{rated}=15$ kW and a nominal voltage of $230\sqrt{2}\,\mathrm{V}$. Therefore, it is suitable for low voltage applications. The control scheme for the DC/AC converter is described in Sec. III-B2.
The parameter values of the DC/DC and DC/AC converter and grid connection filter are summarized in Table I. The overall system is monitored in real-time by a computer, from which the relevant parameters of the controller as well as the auxiliary systems can be modified during runtime using the respective GUI.
Refer to captioncontrol unitDC/DC converter and DC/AC converterpassive components and output filterDC power supply
Figure 6: Overview of a programmable DC/AC converter system for the hardware experiments at the Energy Lab.
IV-A2 Resistive load
For our laboratory experiment, we consider a light-wall shown in Fig. 7. The light-wall consists of 208 light bulbs. Each has a power consumption of $100\,\mathrm{W}$ at nominal voltage of $230\sqrt{2}\,\mathrm{V}$. The light bulbs are evenly distributed over the three phases and can be individually controlled via a programmable logic controller to reach a user-defined load power. Due to the ability of synchronous switching of the light bulbs, variable load steps can be realized using the automation system of the Energy Lab.
IV-A3 Transmission line replica
The transmission line replicas are deployed to reproduce the relevant physical properties of the transmission system [7]. For our experiments, the resistive and inductive or (RL)-type replicas are used. They consist of a resistance set in series with an inductance. The parameter values of the transmission line of our experimental setup are given in Table II.
Figure 7: Resistive load represented by a light-wall.
IV-B Test scenarios
In the remainder, we consider the following test scenarios for our hardware experiments.
  1. 1.
    Scenario I: A single DC/AC converter system in closed-loop with the angular droop control is connected to a resistive load. The goal of this scenario revolves around testing:
    • the black start capabilities, namely the ability to form sinusoidal wave after a major event, e.g., a blackout.
    • the capability to withstand load disturbance, e.g., upon a sudden increase/decrease in the load power.
    • the direct and indirect implementation schemes of the angular droop behavior and drawing a comparison between them.
  2. 2.
    Scenario II: Two identical DC/AC converter systems, each in closed-loop with the angular droop are connected to a common resistive load. In this scenario, we test:
    • the frequency synchronization capabilities of the angular droop control.
    • the power sharing capabilities in dependence of the angular droop control gains.
TABLE I: Technical details of the hardware setup
DC/DC converter and its control
Symbol Definition Range Value in S.I.
$V^{nom}_{dc}$ DC voltage >0 750
$V_{b}$ source voltage >0 600
$C_{dc}$ DC capacitance >0 $3\cdot 10^{-3}$
$R_{b}$ parasitic resistance >0 $1\cdot 10^{-3}$
$L_{b}$ boost inductance >0 $2.36\cdot 10^{-3}$
$k_{P}$ P-voltage gain >0 $0.3$
$k_{I}$ I-voltage gain >0 $12$
$k_{BP}$ P-current gain >0 $10$
$k_{BI}$ I-current gain >0 $200$
DC/AC converter and its control
Symbol Definition Range Value in S.I.
$V^{*}$ AC voltage amplitude >0 $230\sqrt{2}$
$C$ AC filter capacitance >0 $1\cdot 10^{-5}$
$L$ AC filter inductance >0 $2.36\cdot 10^{-3}$
$R$ AC filter resistance >0 $1\cdot 10^{-3}$
$G$ load resistance >0 $58.77$
$A$ modulation amplitude >0 $0.8132$
angular droop control
Symbol Definition Range Value in S.I.
$P^{*}$ active power >0 $2880$
$\omega^{*}$ frequency >0 $2\pi 50$
$\alpha$ input effort gain >0 $2000$
$\gamma$ steady state gain >0 $5\cdot 10^{4}$
$k_{VP}$ P-voltage gain >0 $0.05$
$k_{VI}$ I-voltage gain >0 $0.4$
$k_{IP}$ P-current gain >0 $10$
$k_{II}$ I-current gain >0 $240$
V Scenario I: Single converter to load
V-A Scenario description
This scenario consists of a balanced three-phase DC/AC converter connected to the resistive load and supplied by a DC power supply behind a DC/DC converter as shown in Fig. 8. For the DC/DC converter, we follow vector control to provide a constant voltage level for the DC/AC stage. The DC/AC converter is controlled by the angular droop control either directly or indirectly via the switching modulation signal $\overline{u}$. We refer the reader to Section III-B for the details related to the individual control schemes.
Figure 8: Schematic representation of Scenario I consisting of a balanced three-phase DC/AC converter in closed-loop with angular droop control and connected to a resistive load.
V-B Experimental procedure
We start by configuring the lab automation system [7] according to the hardware configuration in Fig. 8. Thereafter, the DC supply together with the DC/DC converter are enabled and the following experiments are started.
  1. 1.
    Black start: From zero voltage conditions, the DC/AC converter is connected to the grid by closing the respective three-phase circuit breaker. The amplitude of the switching modulation signal in (14) is set to its nominal value given by,
    $A=\frac{2\,V^{d}}{V^{nom}_{dc}},$
    where $V^{d}>0$ is the amplitude of the output voltage $v$ in (15) and $V^{nom}_{dc}>0$ is the nominal DC voltage.
  2. 2.
    Step in the load power: Starting from nominal steady state conditions, we increase the load power at $0.2\,\mathrm{s}$. Moreover, we study suitable tuning of the angular droop control by varying the control gains $\alpha>0$ and $\gamma>0$ and studying their effect on the angle and frequency error transients and their induced steady states.
  3. 3.
    Direct vs. indirect implementation: We compare the black start and step in the load experiments for the DC/AC converter controlled by the angular droop control according to the direct with the indirect implementation described in Section III-B2.
V-C Challenge 1: Discretization of the angle dynamics
The angular droop control described in Section III-B needs to conform with the restrictions posed by the hardware prototype, in particular its hardware implementation in discrete-time. For this, we discretize the closed-loop angle dynamics in Eq. (3) using forward Euler method as follows,
$\theta(s+1)$ $=\theta(s)+T_{s}\,u_{d}(s)+\omega^{*}\mathds{1}_{n}$ (18)
$u_{d}(s)$ $=-\frac{1}{2}R^{-1}\,(\Gamma\,(\theta(s)-\theta^{*}(s))+P(\theta(s))-P^{*})$
where $s\in\mathbb{Z}$ is the time step, $T_{s}>0$ is the sampling period and $\omega^{*}>0$ is the nominal frequency. Next, we define the angle error coordinate $\Delta\theta(s)=\theta(s)-\theta^{*}(s)$ where $\theta^{*}\in^{n}$ denotes the nominal angle satisfying,
$\theta^{*}(s+1)=\theta^{*}(s)+T_{s}\,\omega^{*}.$ (19)
From (18), the discrete-time angle error dynamics are given by,
$\Delta\theta(s+1)$ $=\Delta\theta(s)-\frac{1}{2}T_{s}\,R^{-1}\,(\Gamma\Delta\theta(s)+P(\theta(s))-P^{*}).$ (20)
Observe that in (19), the nominal steady state angle vector $\theta^{*}$ grows infinitely. This causes a loss of precision for the stored variable $\theta(s)$ due to the limit of available bytes for single precision. This observation motivates the following solution: Find a mapping of the angles $\theta^{*}$ from n to the $n-$th dimensional torus $\mathbb{T}^{n}$. Note that the sine function appearing in the implementation of the modulation signal $\overline{u}$ both for the direct and indirect schemes, namely in (14) and (17) is $2\pi-$ periodic in the angle $\theta(s)$. We proceed by limiting the values of the nominal angle $\theta^{*}(s)$ as follows,
[…]
which yields the following absolute angles,
$\theta(s+1)=\theta^{*}(s+1)+\Delta\theta(s+1).$
Here, $\Delta\theta(s+1)$ is given by (20). therefore the angles $\theta(s),\,s\in\mathbb{Z}$ remains within feasible numerical bounds.
V-D Experimental Results
In the following, we present our main observations from the experiments of Scenario I. All unis are in S.I.
  1. 1.
[…]
    Figure 9: Averaged DC voltage $V_{dc}$ in (a), leg inductance current $I_{b}$ in (b), active power $P$ in (c) and AC voltage $v$ in (d) in the black start experiment.
  2. 2.
    Step in the load power:
    Figure 10: Averaged DC voltage $V_{dc}$ in (a), leg inductance current $I_{b}$ in (b), active power $P$ in (c) and AC voltage $v$ in (d) in the load step experiment.
    Figure 11: AC inductance current in the load step experiment
    Following the load step change at $t=0.2$s, the active power overshoots to approximately $1.2\cdot 10^{4}$ and settles at $P^{s}\approx 3800$ at steady state from Fig. 10. This is induced by large rush currents accompanying the change in load power as shown in Figure 11. The DC voltage $V_{dc}$ returns to its nominal value after $0.1\,\mathrm{s}$ due to the integral control action of the inner current loop in (11), whereby the leg inductance current reaches a newly induced steady state $I_{b}^{s}\approx 7$ following the load step. The three-phase voltage amplitude drops slightly during the event and stays around its nominal value at steady state.
    Figure 15 shows a drop in the AC frequency during the load power change, however the frequency returns to its nominal steady state despite the increase in load power. Zero steady state frequency error is a characteristic property of the angular droop control. See also Section II-B.
    In the following, we determine a suitable tuning of the angular droop control by studying the influence of the control gains $\alpha>0$ and $\gamma>0$ on the frequency and angle of the modulation signal following a step in the load at $t=0.2\,$s. As depicted in Fig. 12, an increase in $\alpha$ leads to a higher penalty on the angle and power deviations from their steady state. This translates into a larger rate of change of frequency and smaller nadir. This is in accordance with the observation that, the choice of the gain $\alpha$ affects the transient behavior of the AC frequency $\omega$ and angle $\theta$ in (6). To avoid an overshoot and limit the frequency deviation to an acceptable deviation of $0.8$ Hz [33], we fix the value $\alpha=2000$. Fig. 13 shows that, for decreasing values of $\gamma$, the drop in the angle error at steady state or power-to-angle droop behavior is more significant following (5). This results in an induced steady state angle $\theta^{s}$ given by
    $\theta^{s}=\theta^{*}+\frac{1}{\gamma}(P^{*}-P^{s}),$ (22)
    where $P^{s}$ is the load power at induced steady state following a step in the load. For different $\gamma$ values the rate of change of the angles during the transient leads to frequency behavior described by (6). This empirically confirms that $\gamma$ affects both the transients of the AC frequency $\omega$ and steady state behavior of the angle $\theta$. Since the choice of $\gamma\geq 5\cdot 10^{5}$ leads to relatively small power-to-angle droop behavior, we select $\gamma=5\cdot 10^{4}$.
[…]
  3. 3.
    Comparison between direct and indirect control: For the control gains $\alpha=2000$ and $\gamma=5\cdot 10^{4}$, Fig. 14 and 15 compare the frequency and angle errors resulting from the direct and indirect implementation for the black start and the load step experiment, respectively. Also in the indirect implementation, the frequency is restored to its nominal value within approximately $0.7\,\mathrm{s}$ and the angles converge to an induced steady state angle $\theta^{s}$ as in Eq. (22). In fact, the major difference between direct and indirect control is in the way the control law (14) relates to the modulation signal representing the main input to the DC/AC converter. In the direct implementation, we assign a sinusoidal wave whose angle is directly determined by the angular droop control. The indirect implementation encodes a nested voltage and current control loops relying on tracking a given AC voltage reference, whose angle is described by the angular droop control. This results in different modulation signals $\bar{u}$, where the control effort solely dependent on the gains $\alpha$ and $\gamma$ in the direct scheme. For the indirect scheme, the control effort is also dependent on the choice of the inner (current) and outer (voltage) control loops, see Table I. For our particular control gain choice, this results in a slower convergence rate to a steady state compared to the direct scheme. This is visible from the duty cycles $d$ in Fig. 16, defined by (10). The duty cycle increases linearly in the direct whereas sub-linearly in the indirect implementation scheme.
    Due to its simpler and more intuitive tuning, we adopt the direct implementation of the angular droop control with $\alpha=2000$ and $\gamma=5\cdot 10^{4}$ in the remainder of our hardware experiments.
    Figure 14: Frequency (a) and angle errors (b) for the black start experiment.
    Figure 15: Frequency (a) and angle errors (b) in the load step experiment.
    Figure 16: Duty cycles of the indirect (a) and direct control (b) schemes in the black start experiment.
VI Scenario II: Two identical converters to a common load
VI-A Scenario description
This scenario consists of two identical three-phase balanced DC/AC converters supplied by two independent DC sources behind DC/DC converters and connected to a common resistive load as shown in Fig. 17. The control scheme of the boost converters follow the vector control discussed in Section III-B1. The DC/AC converters are in closed-loop with the angular droop control according to the direct implementation described in Section III-B2. Table II summarizes the parameter values. To satisfy the modeling assumptions in Section II-A, we set the ratio $X/R_{\ell}\approx 11$ with the reactance $X=\omega^{*}L_{\ell}$, corresponding to highly inductive power network [34, 26].
TABLE II: Parameters values of the transmission line.
Symbol Definition Range Value in S.I.
$R_{l}$ line resistance >0 $20\cdot 10^{-3}$
$L_{l}$ line inductance >0 $700\cdot 10^{-6}$
Figure 17: Schematic representation of Scenario II consisting in two identical converter systems, each in closed-loop with angular droop control, connected to a common resistive load via highly inductive transmission lines. The line resistance is negligible.
VI-B Experimental procedure
We start by connecting the DC/AC Converter I following the steps described in Section V-B) to the resistive load. From this initial setup, we conduct the following experiments:
  1. 1.
    Frequency synchronization: we connect the DC/AC Converter II at $t=0$ to the initial setup consisting in Converter I connected to resistive load by closing the connection relay and enabling its modulation.
  2. 2.
    Power sharing: Starting from a frequency synchronous two-converter system, we study in experiments the power sharing capabilities of the angular droop control by suitable tuning of the power-to-angle droop gains.
VI-C Challenge 2: Clock drift in angular droop controllers
VI-C1 Analysis
Angular droop control is susceptible to clock drifts. This is well-documented in the power system literature [26, 35, 36, 37] and can be explained as follows. Let $t_{k}>0$ denote the local time at the $k$-th DC/AC converter with respect to a global reference time $t$ as follows,
$t_{k}:=(1+\epsilon_{k})t,$ (23)
where $\epsilon_{k}>0$ is the time-invariant drift of the local clock with respect to the reference clock. This drift arises in the absence of a master clock. Starting from a nominal global frequency $\dot{\theta}^{s}_{k}=\omega^{*}$, the angle obtained from the integration (assuming zero initial conditions) is affected by the clock drift because
$\theta^{s}_{k}(t)=\omega^{*}\,t_{k}=\omega^{*}\,(1+\epsilon_{k})\,t=\omega_{k}\,t,$ (24)
where $\omega_{k}:=\omega^{*}\,(1+\epsilon_{k})$ is local frequency at the $k-$th converter. Under the assumption of a highly inductive, Kron-reduced [38] power network, the active power at the output of the $k$-th converter at steady state is given by [2]
$P^{s}_{k}$ $=\frac{V_{k}\,V_{j}}{X_{kj}}\sin(\theta^{s}_{k}-\theta^{s}_{j})=\frac{V_{k}\,V_{j}}{X_{kj}}\sin((\omega_{k}-\omega_{j})t)$
$=\frac{V_{k}\,V_{j}}{X_{kj}}\sin(\omega^{*}\,t\,(\epsilon_{k}-\epsilon_{j})),$
where $\theta^{s}_{j}(t):=\omega_{j}\,t$ is the steady state angle, $\omega_{j}$ is the local frequency at the $j-$th converter and $X_{jk}$ is the reactance between converters $k$ and $j$ following the Kron-reduction. It can be deduced that when local clock drifts are not compensated for, the injected active power $P^{s}_{k}$ drifts apart from its nominal value. Since the angular droop control law involves an integration, the closed-loop dynamics (3) are not robust to local clock drifts [37] and suitable solution needs to be developed for the hardware implementation.
VI-C2 Proposed solution
In our lab experiments, we utilize Imperix distributed modulation allowing for synchronized PWM signals generation across multiple digital controllers [39]. This allows multiple controllers to behave as if all PWM signals were generated from a single centralized modulator. The distributed modulators must provide the same time resolution and accuracy as modulators belonging to a unique hardware. This technology consists in the distribution of a common high-frequency clock (or master clock) across the entire control network through a direct optical fiber connection between the clocks of Converter I and II [40]. Therefore, the distributed devices belong to same clock domain, which eliminates the clock drift for the integration actions and allows for distributed modulation with an accuracy of ±$\pm 2\,\mathrm{ns}$.
VI-D Experimental results
Figure 18: Frequency (a) and angle error (b) for $k=1,2$ and angle differences (c), following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ for $k=1,2$. Here $\theta^{*}_{1}(t)=\omega^{*}t$ for $t<0$ and $\theta^{*}_{2}(t)=\omega^{*}t+\theta_{1}(0)-0.042$ for $t\geq 0$, where $\theta_{1}(0)$ is the modulation angle of Converter I at the time of interconnection.
Figure 19: Representation of Scenario II as two voltage sources with switching voltage amplitude $V_{k}$ and modulation angle $\theta_{k}$ for $k=1,2$ supplying a common resistive load under the assumption of a high $X/R_{l}$ ratio with line reactance $X=\omega^{*}L_{l}$ and constant voltage amplitudes.
VI-D1 Frequency synchronization
Figure 18 depicts the frequency and angle errors following the connection of Converter I to Converter II. Before the interconnection, the modulation angle of Converter I is initialized at zero with $\theta_{1}^{*}(t)=\omega^{*}t$, for $t<0$. The nominal angle of Converter II is given by $\theta^{*}_{2}(t)=\omega^{*}t+\theta_{2}(0),\,t\geq 0$, where $\theta_{2}(0)$ is the initial angle of Converter II at the time of interconnection $t=0$. The choice of the initial angle $\theta^{*}_{2}(0)$ can be determined as follows. The two-converter system represented in Fig. 17 can be reduced to two sources connected to one load as depicted in Fig. 19. The active power at the $k$-th DC/AC converter at steady state is given by [2]
$P^{s}_{k}=\frac{V_{k}\,V_{0}}{X_{k0}}\sin{(\theta^{s}_{k}-\theta^{s}_{0})}\,,\,k=\{1,2\}$ (25)
where $V_{0}\angle\theta^{s}_{0}$ is the phasor at the common node connecting the two converters and $V_{k}\angle\theta^{s}_{k}$ that of the switching voltage of the $k-$th DC/AC converter with $V_{k}=\lVert\frac{1}{2}u_{k}V_{dc}\rVert$ and $X_{k0}=\omega^{*}L_{l}>~{}0$ is the reactance of the inductive line impedance. Thus, the active power at Converter II is given by $P^{s}_{2}=\frac{V_{2}\,V_{0}}{X_{20}}\sin{(\theta^{s}_{2}-\theta^{s}_{0})}$. Setting $P^{s}_{2}=P^{*}_{2}=0$ leads to $\theta_{2}(0)=\theta_{0}(0)$. From $P^{s}_{1}=\frac{V_{1}\,V_{0}}{X_{10}}\sin{(\theta^{s}_{1}-\theta^{s}_{0})}=P^{*}$, we obtain
$\theta_{1}(0)=\theta_{0}(0)+\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}}),$ (26)
where $\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})=0.006$, $\theta_{1}(0)$ is the initial angle of Converter I and $\theta_{0}(0)$ is the initial angle at node 0 depicted in Fig. 19 obtained from a Phased-Locked-Loop (PLL) scheme [41] at the time of interconnection. Our experimental results in Fig. 18 show that the modulation angle differences at steady state, i.e., $\theta^{s}_{1}-\theta^{s}_{2}=0.002$. This can be inferred from the negligible but non-zero line resistance $R_{l}>0$ leading to inevitable power mismatches, also visible in Fig. 20, as follows
$\theta_{1}^{s}-\theta_{2}^{s}$ $=\theta_{1}^{*}-\theta_{2}^{*}+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=\theta_{1}(0)-\theta_{0}(0)+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=\theta_{0}(0)+\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})-\theta_{0}(0)+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=\arcsin(\frac{P^{*}X_{10}}{V_{0}V_{1}})+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2})$
$=0.006+\frac{1}{\gamma}(\delta P_{1}+\delta P_{2}),$
where $\delta P_{1}+\delta P_{2}=-200$ and $\delta P_{1}=P^{*}_{1}-P^{s}_{1}$ and $\delta P_{2}=P^{*}_{2}-{P}^{s}_{2}$. Thereby, the steady state angle differences remain within $[-\pi/2,\pi/2]$ (rad) and the security constraint in Assumption 1 is satisfied. Under Assumption 1, the two converters synchronize at nominal frequency $\omega^{*}$ within $0.25\,s$. Both converters’ angles converge to their frequency synchronous steady states. This corresponds to the active power of Converter I and II, $P^{s}_{1},P^{s}_{2}$ reaching a nearby nominal value as seen in Figure 20. We note thereby that $P^{*}_{1}+P^{*}_{2}=P^{*}$ where $P^{*}$ is the total active power drawn by the load resistance. Furthermore, the converters exchange reactive power $\tilde{Q}^{s}_{1}<0$ and $\tilde{Q}^{s}_{2}>0$, where $\tilde{Q}^{s}_{1}+\tilde{Q}^{s}_{2}=0$ at all times $t\geq 0$. This can be explained by the unavoidable mismatches of the voltage amplitudes (see also Fig. 21), which leads to non-zero reactive power. In particular, in a highly inductive, Kron-reduced [38] power network, the reactive power is expressed at the $k-$th converter by [42]
$\tilde{Q}^{s}_{k}=\frac{V_{k}}{X_{kj}}(V_{k}-V_{j}\cos(\theta_{k}-\theta_{j})),$ (27)
where $k\neq j$ and $k=\{1,2\}$. Under the small signal approximation, i.e., $\theta_{k}-\theta_{j}\approx 0$, we obtain
$\tilde{Q}^{s}_{k}\approx\frac{V_{k}}{X_{kj}}(V_{k}-V_{j}).$ (28)
Hence, $\tilde{Q}^{s}_{k}\neq 0$ for $k=\{1,2\}$ due to $V_{k}\neq V_{j}$. Finally, the phase portrait of the output voltages $v_{1}$ and $v_{2}$ in Fig. 21 represents a limit cycle of an approximate radius of $V^{d}$ in the phase plane which shows once again the frequency synchronization in AC voltages nearby a desired voltage amplitude $V^{d}$. As a conclusion, our experiment validates the local asymptotic stability result shown in [20].
Figure 20: Active (a) and reactive (b) power $P^{s}_{k}=v_{k}^{\top}\,i_{o,k}$ and $Q^{s}_{k}=v_{k}^{\top}\,\mathbf{J}\,i_{o,k}$ with $P_{1}^{*}=P^{*},P^{*}_{2}=0$ with $k=1,2$ for the frequency synchronization experiment. Converter I and II in closed loop with direct angular droop with $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ and $k=1,2$.
Figure 21: Phase portrait of the periodic orbit of the three-phase output voltages $v_{1}$ and $v_{2}$ in $abc$- frame, respectively in closed-loop with the direct angular droop control for $\gamma_{k}=5\cdot 10^{4}$ and $\alpha_{k}=2000$ and $k=1,2$ during the frequency synchronization experiment.
VI-D2 Power sharing
Figure 22: Frequency (a), angle error (b) for $k=1,2$ and angle differences (c), following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=500$ and $\alpha_{k}=2000$ for $k=1,2$ and $r=1$ in the power sharing experiment.
Figure 23: Active power $P_{k}=v_{k}^{\top}i_{o,k}$ in (a) and reactive power $Q_{k}=v_{k}^{\top}\mathbf{J}i_{o,k}$ in (b) following the connection of Converter II to Converter I at $t=0$, both in closed-loop with the direct angular droop control for $\gamma_{k}=500,\alpha_{k}=2000$ and the power ratio $r=1$ in (33) for $k=1,2$ in the power sharing experiment, $P_{1}^{*}=P^{*}_{2}=\frac{P^{*}}{2}$.
To achieve power sharing among the Converters I and II, we first determine a suitable tuning for the power-to-angle droop gains $\gamma_{k},\,k=1,2$ and the nominal power ratio $P_{1}^{*}/P_{2}^{*}$, by conducting the following analysis inspired by [26]. To keep the analysis tractable, we assume in the remainder an inductive power network characterized by high ${X}/{R_{l}}$ ratio (see Table II), constant voltage magnitudes and neglect the output filter at each converter.
Choice of $(\gamma_{k},P^{*}_{k}),\,k=1,2$
Under the small signal approximation, (25) can be rewritten as
$(\theta^{s}_{k}-\theta^{s}_{0})\approx\frac{X_{k0}}{V_{k}\,V_{0}}P^{s}_{k}\,.$ (29)
At steady state, the angular droop control law is given by (5), where
$\theta^{s}_{k}$ $=\theta_{k}^{*}+\frac{1}{\gamma_{k}}(P^{*}_{k}-P^{s}_{k}),\quad k={1,2}.$ (30)
Here $\theta_{k}^{s}\in$ and $P_{k}^{s}>0$ are the induced kron-reduced steady state angle and active power at the $k-$th converter.
By letting $\theta^{s}_{1}-\theta^{s}_{2}=\theta^{s}_{1}-\theta^{s}_{0}+\theta^{s}_{0}-\theta^{s}_{2}$, we obtain,
$\theta^{*}_{1}-\theta^{*}_{2}-\frac{1}{\gamma_{1}}(P^{s}_{1}-P^{*}_{1})+\frac{1}{\gamma_{2}}(P^{s}_{2}-P^{*}_{2})=\frac{X_{10}}{V_{1}\,V_{0}}P^{s}_{1}-\frac{X_{20}}{V_{2}\,V_{0}}P^{s}_{2}.$ (31)
By reordering the terms in (31), we arrive at
$\theta^{*}_{1}$ $-\theta^{*}_{2}+\frac{1}{\gamma_{1}}P_{1}^{*}-\frac{1}{\gamma_{2}}P_{2}^{*}$
$=\bigg{(}\frac{1}{\gamma_{1}}+\frac{X_{10}}{V_{1}\,V_{0}}\bigg{)}P^{s}_{1}-\bigg{(}\frac{1}{\gamma_{2}}+\frac{X_{20}}{V_{2}\,V_{0}}\bigg{)}P^{s}_{2}.$
From (26), we have that $\theta^{*}_{1}(t)-\theta^{*}_{2}(t)=\theta_{1}(0)-\theta_{0}(0)=0.006$. Therefore, if we select the power-to-angle-droop gain $\gamma_{k}>0$ such that,
$\gamma_{k}\ll\frac{V_{k}V_{0}}{X_{k0}},\quad k=1,2$ (32)
holds, it yields that
$\frac{1}{\gamma_{1}}P_{1}^{*}-\frac{1}{\gamma_{2}}P_{2}^{*}\approx\frac{1}{\gamma_{1}}P^{s}_{1}-\frac{1}{\gamma_{2}}P^{s}_{2},$
and we deduce that for an active power ratio defined as,
$r:=\frac{P^{*}_{1}}{P^{*}_{2}}=\frac{\gamma_{1}}{\gamma_{2}},$ (33)
it holds that,
$\frac{P^{s}_{1}}{P^{s}_{2}}\approx r,$ (34)
and the power sharing amongst the two converters is guaranteed. Finally, for our experimental setup, the condition (32) can be rewritten as
$\gamma_{k}\ll 4.8\cdot 10^{5},$ (35)
with $V_{1}=V_{2}=V^{d},V_{0}\approx V^{d}$ and $X_{10}=X_{20}=\omega^{*}L_{l}$. To achieve (35), we select $\gamma_{1}=\gamma_{2}=500$ with $r=1$ throughout the power sharing experiment.
Discussion
Figures 22 and 23 show the experimental results of the power sharing experiment. Assumption 1 is here again satisfied and that is seen from Figure 22, where $\lVert\theta^{s}_{1}-\theta^{s}_{2}\rVert<0.1$. For $\gamma_{1}=\gamma_{2}=500$, power sharing is guaranteed at steady state, where $P^{s}_{k}\approx P^{*}/2$ for $k=1,2$. This corresponds to zero reactive power at steady state. As expected, the two converters synchronize in frequency and their angles converge respectively to frequency synchronous steady states within 0.50.50.5\,0.5s. The trade-off between active power sharing and the induced steady state modulation angle is visible when comparing Figure 18 and 22. Thus, our experiment validates the power sharing capabilities of the angular-droop controlled DC/AC converter system, one of the most important plug and play properties for converter control design in power networks.
VII Conclusion
We validated the grid-forming capabilities of the angular droop control through two test scenarios embedded within the hardware experimentation in the Smart Energy System Control Laboratory at KIT. In both scenarios, the DC/AC converter in closed-loop with the angular droop control recovers nominal operation following a blackout, provides active power, also upon a sudden change in the load, achieves frequency synchronization and power sharing in a network setting upon suitable tuning. Our future work aims at addressing the compatibility of the angular with frequency-droop control and therefore the interoperability with the remainder of the grid components. Another venue is concerned with voltage control and large-scale hardware experimentation.